sam took his family to the zoo. an adult's ticket is two times the cost of a child's ticket. the total cost for two adult's tickets and three children's tickets was $28. how much do the tickets cost?
Answer by mananth(16946)

child's ticket = x
adult's ticket = 2x
..
child tickets = 3
adults tickets = 2
...
3x+2*2x=28
3x+4x=28
7x=28
/7
x= 28/7
x= $4 child ticket
adult ticket = 2x = 2*4 = 8